# Large Twist Angle dependent Ultrafast Transient Dynamics and Raman studies on MoSe2/WSe2 van der Waals Heterostructures

### v1

@@ 1 Optical pump - optical probe studies
We offer a qualitative understanding of our results. We observed that $\tau_{2}$ is maximum at the twist angle of 30, i.e. the interlayer excitons take the longest time to recombine. This is due to the interlayer distance being maximum at the twist angle of 30 in TMDs HSs[29, 30]. Equally interesting is the observation of two minima of $\tau_{2}$ observed in the range of 20-25 and 35-38. These two ranges of twist angles cover the commensurate angles of $\theta$ = 21.8 and 38.2 for MoSe2/WSe2 HS, where interlayer excitons can access an additional channel to recombine[26, 31]. Seyler et al. have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at $\pm$±K points do not align in the first Brillouin zone. However, at the commensurate angle of 21.8, the valence band and conduction band align at $\pm$±K points in the second Brillouin zone, as shown in Figure 4, adapted from ref. [26]. These excitons can recombine radiatively, and the momentum mismatch (due to non-zero twist) is compensated by the reciprocal lattice of two layers; the process known as Umklapp recombination [31, 26]. We attribute this Umklapp recombination in the vicinity of commensurate angles of 21.8 and 38.2 as a cause for the anomalous minima of $\tau_{2}$. The two different fluences support the evidence of anomaly of relaxation dynamics in the vicinity of commensurate twist angles. Although the fraction of intralayer excitons increases with an increase in fluence (Figure 3d), the absolute numbers of interlayer and intralayer excitons increase with an increase in the fluence across all twist angles. This increment in the number of interlayer excitons results in longer recombination time as fluence increases (lower panel of Figure 3d). The fluence dependence at 20, near the commensurate angles, displays that with the increase in fluence, there is no significant increase in the fraction of intralayer to interlayer excitons (A1/(1-A1)), suggesting that both types of excitons increase proportionally. It is important to note that the $\tau_{2}$ value remains approximately constant at 20 with increasing fluence, even after a significant increase in the number of interlayer excitons compared to other twist angles. In conclusion, additional recombination channels for interlayer excitons at commensurate angles enhance their saturation density and keep the recombination time ($\tau_{2}$) independent of fluence within the reported range.
2 Twist angle dependent Raman studies in MoSe2/WSe2 HS
The blue and red shifts of the A1g modes in WS2 and MoS2 within their twisted HSs [30] motivated us to perform Raman spectroscopy on HS MoSe2/WSe2 for various twist angles. The group theory for the individual layer provides the irreducible decomposition of Raman active vibration modes A1g(R), E2g(R), and E1g (IR+R) for MX2 monolayer [54]. For the HS, we obtain the following vibration modes as shown in Figure 1c: M1(MoSe2, A1g): 238.5 cm-1, M2(WSe2, A1g+E2g): 249 cm-1, M3(WSe2, 2LA): 258.5 cm-1, M4(MoSe2, E2g): 286 cm-1, M5(WSe2, A′′${}_{2}^{\prime\prime}$): 305 cm-1, M6(MoSe2, A′′${}_{2}^{\prime\prime}$): 353 cm-1, M7: 373 cm-1 and M8: 394 cm-1 [55, 56]. From the perspective of the charge transfer among two vertically stacked monolayers, we concentrate on the out-of-plane modes (A1g) of the two monolayers MoSe2 and WSe2 stacked together, namely M1 (MoSe2) and M2 (WSe2).
Figure 5a shows the normalized intensities of M1, M2, and M3 with different twist angles. The black dashed lines represent the shift of the peak position of modes M1 or M2 at different twist angles. This systematic behavior of M1 and M2 can be seen in Figure 5b, where the dashed red lines (a guide to the eye) show the opposite slopes for M1 and M2 modes for four different ranges of the twist angles. We observed the softening of M1 and hardening of M2 mode for any non-zero twist as compared to the aligned geometry (depicted by blue circles for zero twist angle in Figure 5a). Mode M2 starts stiffening as the twist angle is varied till 17, followed by softening till 32 and then stiffening again till 39 which is again followed by softening. In a counter behavior, the M1 mode softens as the twist angle is varied till 17, followed by stiffening till 32 and then softening till 39, which is again followed by stiffening. The softening and stiffening of the A1g mode (both M1 and M2) depend upon the carrier doping as well as the interlayer coupling [49, 57, 58]. In transition metal dichalcogenide MoS2, increasing the number of layers stiffens the A1g mode and softens the E2g mode due to increased dielectric screening of Coulomb interactions [59, 60]. It has been reported that the unidirectional transfer of electrons from monolayer WS2 to monolayer MoS2 leads to higher carrier density in MoS2 and lower carrier density in WS2, which can soften or stiffen the respective A1g modes[30]. For the similar staggered bands configuration of HS MoSe2/WSe2, unidirectional electron flow from WSe2 to MoSe2 increases (reduces) the carrier density in MoSe2 (WSe2) and leads to the softening (stiffening) of M1 (M2) modes. To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 5c. A guide to the eye (solid red curve) shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8 and 38.2. For perfectly aligned HS ($\theta$=0), the electron transfer and the interlayer coupling lead to the softening of A1g of both M1 and M2 modes, if we compare these modes in monolayers of MoSe2 and WSe2 (Figure 5a). However, with a non-zero twist angle, variation in the charge transfer and the interlayer coupling tunes the softening and the corresponding stiffening of M1 and M2 modes and shows an anomalous maximum shift at the angle of 19 and 39. This can be understood as the maximum electron transfer and more dielectric screening of Coulomb interaction between electron and hole for the MoSe2/WSe2 HS at these specific angles, which are in the vicinity of commensurate angles. This anomalous behavior led us to observe the evolution of other modes of HS with twist angle.
The disappearance of M3 (2LA mode of WSe2) at the twist angles of 19 and 39 as depicted by small arrows in Figure 5a (panel iii) is observed. Interestingly, this observation for M3 mode happens in the vicinity of commensurate angles 21.8 &\&& 38.2 which needs further investigation. Thus, the characteristics of M1 and M2 modes provide us with strong evidence of maximum equilibrium charge transfer in the MoSe2/WSe2 HS at commensurate angles.
Optical pump-optical probe spectroscopy and Raman spectroscopy reveal that charge transfer between MoSe2 and WSe2 layers in the heterobilayer is maximum near commensurate angles of 21.8 and 38.2. This increased charge transfer leads to more screening of Coulomb interactions between electrons and holes, resulting in a shorter lifetime for interlayer excitons. This decrease in $\tau_{2}$ is also accompanied by the additional recombination channel provided by the Umklapp process.
@@ Conclusions
In summary, we have reported the optical pump optical probe spectroscopy of MoSe2/WSe2 HS with different twist angles. We observed a significant decrease in interlayer exciton recombination time in the proximity of commensurate angles of 21.8 and 38.2. We interpret the result as a consequence of the availability of new radiative recombination channels as a result of Umklapp recombination. Raman spectra show the systematic softening and stiffening of the A1g modes M1 and M2 with the non-zero twist angle, a consequence of unidirectional charge transfer in TMDs with type-II band configuration. The maximum shift in modes M1 and M2 has been observed again in the vicinity of commensurate angles. Our results can be helpful in designing the applications of twisted vdW HS in optoelectronic applications.
Conflicts of interest
There are no conflicts of interest to declare.
@@ Acknowledgments
AKS thanks the Department of Science and Technology for its financial support under the National Science Chair Professorship. VA acknowledges CSIR for the research fellowship. PKN acknowledges the MHRD STARS research grant [STARS/APR2019/396].
Figure 1: (a) Microscopic image of MoSe2/WSe2 HS depicting WSe2 as a bottom layer (big triangles) with dull contrast and MoSe2 as a top layer (smaller triangles) with brighter contrast, (b) Angle evaluation between two monolayers with the blue(red) triangle as a boundary of WSe2(MoSe2) monolayers, (c) The Raman spectrum for HS at a $\theta$ = 49. M1(M2) represents the A1g mode of MoSe2(WSe2), The respective modes are as follows: M1(238.5 cm-1), M2(249 cm-1), M3(258.5 cm-1), M4(286 cm-1), M5(305 cm-1), M6(353 cm-1), M7(373 cm-1), M8(394 cm-1).
Figure 2: Type II band structure of HS MoSe2/WSe2, illustrating the charge transfer and the formation of the intralayer exciton (IaX) and interlayer exciton (IeX) at different time intervals after photoexcitation.
Figure 3: (a) Time evolution of normalized $\Delta$R/R for MoSe2/WSe2 HS at a $\theta$=20, inset: the data up to 30 ps, the blue dots represent the data points and the red solid curve represents the fitting model given by Eq. 1, (b) The cumulative plot of normalized $\Delta$R(t)/R emphasizes the twist angle ($\theta$) dependence of the transient dynamics at a fluence of 2.0 mJ/cm2, (c) i. the fractional amplitude and ii. recombination time of interlayer excitons of HS at different twist angles for two different fluences (blue circle: $\phi_{1}$=2.0 mJ/cm2, red triangle: $\phi_{2}$=2.8 mJ/cm2). The solid blue line is a guide to the eye, (d) The fluence dependence of the fraction of intralayer to interlayer exciton density, and $\tau_{2}$ for twist angles of 4, 20, and 28.
@@
Figure 5: (a) The graphs depict i. M1, ii. M2 and iii. M3 modes at different twist angles and constituent monolayers. The corresponding stiffening and softening of M2 and M1, respectively, with the twist angle are indicated by the black dashed line, M3 mode disappears at $\theta$=18 and 39 as indicated by solid arrows in iii. (b) The frequency of M1 and M2 modes is plotted as a function of twist angle ($\theta$), the dashed red lines represent the opposite trend of phonon frequencies of M1 and M2 modes with $\theta$ in the different twist regions, (c) The difference of frequencies of modes M1 and M2 is presented with twist angles. The difference is maximum in the proximity of commensurate angles of 21.8 and 38.2.

### v2

@@ Introduction
We offer a qualitative understanding of our results. We observed that $\tau_{2}$ is maximum at the twist angle of $\sim$30 (Figure 5b), i.e. the interlayer localized carriers take the longest time to recombine. This is due to the interlayer distance being maximum at the twist angle of 30 in TMDs HSs[28, 29]. Equally interesting is the observation of two minima of $\tau_{2}$ observed near 20 and 38. These twist angles are close to the commensurate angles of $\theta$ = 21.8 and 38.2 for MoSe2/WSe2 HS, where interlayer localized charge carriers or/and excitons can access an additional channel to recombine[27, 30]. Seyler et al. have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at $\pm$±K points do not align in the first Brillouin zone. However, at the commensurate angle of 21.8, the valence band and conduction band align at $\pm$±K points in the second Brillouin zone, as shown in Figure 6, adapted from ref. [27]. At these specific angles, the interlayer charge carriers can recombine radiatively, and the momentum mismatch (due to non-zero twist angle) is compensated by the reciprocal lattice of two layers; the process known as Umklapp recombination [30, 27]. We attribute this Umklapp recombination in the vicinity of commensurate angles as a cause for the two minima observed for $\tau_{2}$ (Figure 5b). The two different fluences support the evidence of faster relaxation dynamics in the vicinity of commensurate twist angles. The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a minimum in A3 near these commensurate angles (Figure 5c). To further investigate the minima for $\tau_{2}$ near the commensurate angles, the relaxation dynamics is studied as a function of excitation density for different twist angles.
The fluence dependence of the relaxation dynamics of interlayer charge carriers (Figure 5d) shows an increase in interlayer charge carriers with increasing fluence, which saturates beyond 2.0 mJ/cm2 indicating the threshold of charge transfer in HS MoSe2/WSe2 at fluences above 2.0 mJ/cm2. It is important to note that $\tau_{2}$ is least fluence dependent at 20 and 38 in comparison to other twist angles, despite a significant increase in the number of interlayer charge carriers. Intralayer charge carriers do not exhibit this distinct behavior near commensurate angles, as shown by the fluence dependence of amplitude A1 and $\tau_{1}$ in supplementary Figure S3. In conclusion, the additional recombination channels for interlayer charge carriers at commensurate angles demonstrate the fluence independence of recombination time ($\tau_{2}$) within the reported range up to 2.8 mJ/cm2.
@@ Conclusions
In summary, we have reported the optical pump optical probe spectroscopy of MoSe2/WSe2 HS with different twist angles at high excitation densities above the Mott transition. We observed a significant decrease in interlayer exciton recombination time in the proximity of commensurate angles of 21.8 and 38.2. We interpret the result as a consequence of the availability of new radiative recombination channels due to Umklapp recombination. The least fluence dependence of interlayer charge carriers recombination time near commensurate angles further supports the presence of additional recombination channels. Our results can be helpful in designing the applications of twisted vdW HS in optoelectronic applications.
@@ Acknowledgments
AKS thanks the Department of Science and Technology, Government of India, for its financial support under the National Science Chair Professorship and Nanomission (Grant No. DST/NM/TUE/QM-5/2019). DVSM also acknowledges the DST for the Nanomission grant. VA acknowledges CSIR for the research fellowship. PKN acknowledges the MHRD STARS research grant [STARS/APR2019/396] and the support from the Institute of Eminence scheme at IIT-Madras through the 2D Materials Research and Innovation Group.
Figure 1: (a) Microscopic image of MoSe2/WSe2 HS depicting WSe2 as a bottom layer (big triangles) with dull contrast and MoSe2 as a top layer (smaller triangles) with brighter contrast, (b) Angle evaluation between two monolayers with the blue(red) triangle as a boundary of WSe2(MoSe2) monolayers.
Figure 2: (a) The Raman spectrum for HS at a $\theta$ = 49. M1(M2) represents the A1g mode of MoSe2(WSe2), The respective modes are as follows: M1(238.5 cm-1), M2(249 cm-1), M3(258.5 cm-1), M4(286 cm-1), M5(305 cm-1), M6(353 cm-1), M7(373 cm-1), M8(394 cm-1). (b) The frequency of M1 and M2 modes is plotted as a function of twist angle ($\theta$), the dashed red lines represent the opposite trend of phonon frequencies of M1 and M2 modes with $\theta$ in the different twist regions, (c) The difference of frequencies of modes M1 and M2 is presented with twist angles. The difference is maximum in the proximity of commensurate angles of 21.8 and 38.2.
Figure 3: (a) Time evolution of $\Delta$R/R for ML MoSe2 (blue circles), ML WSe2 (red circles), HS MoSe2/WSe2 at $\theta$=20 (pink circles), $\theta$=35 (black circles). (b) The differential reflectivity $\Delta$R/R for HS MoSe2/WSe2 at twist angle of $\theta$=20 (blue circles) is fitted using Eq.(1) (red solid line). The inset shows the fit of the experimental data up to 30 ps.
Figure 4: Type II band structure of HS MoSe2/WSe2, (a) Photoexcitation: Electrons are transferred to the conduction band of MoSe2 and WSe2, Charge transfer: electrons (holes) transfer from conduction (valence) band of WSe2 (MoSe2) to MoSe2 (WSe2), Carriers Localization: interlayer charge carriers localized to distinct individual layers. (b) intralayer e-h recombination (A1, $\tau_{1}$): electrons and holes within the same layer recombine radiatively, Interlayer e-h recombination (A2, $\tau_{2}$): Electron and holes in different layers recombine radiatively, Interlayer Auger recombination (A3): Electrons and holes in distinct layers recombine non-radiatively, transferring energy to a third particle, which is either an electron or a hole.
Figure 5: Twist-angle dependence of (a) amplitude (A1) and recombination time ($\tau_{1}$), (b) amplitude (A2) and recombination time ($\tau_{2}$), (c) amplitude (A3) at a fluence of 2.0 mJ/cm2. The solid lines in (b) and (c) are guides to the eyes. (d) Fluence dependence of amplitude (A2) and recombination time ($\tau_{2}$) for different twist angles. The dashed and dotted (red and blue) lines are guides to the eyes for the data corresponding to 1 and 20